OpenStudy (anonymous):

If f(x) = x2 + 2x + 3, what is the average rate of change of f(x) over the interval [-4, 6]?

OpenStudy (anonymous):

What do you think it is?

OpenStudy (anonymous):

I think it's 4?

OpenStudy (anonymous):

(A).51 (B).40 (C). 31(D). 20 (E).4

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (trojanpoem):

Yep ?

OpenStudy (trojanpoem):

Average = F(6)- F(-4) / 6 + 4 = 27 + 13 / 6 + 4 = 4

OpenStudy (trojanpoem):

(e)

OpenStudy (anonymous):

The graph shows a line joining two complex numbers. What are the two complex numbers?

OpenStudy (anonymous):

Should I post the multiple choice answers?

OpenStudy (trojanpoem):

Ok.

OpenStudy (trojanpoem):

0 + 3 i 3 - 3 i

OpenStudy (trojanpoem):

Hey ?

OpenStudy (anonymous):

A. -3 and 3-3i B. -3i and 3+3i C. 3 and 3i-3 D. 3i and 3-3i E. 3i and 3i-3

OpenStudy (trojanpoem):

3i and 3- 3 i

OpenStudy (trojanpoem):

I sent the answer to you 2 years ago.

OpenStudy (anonymous):

Computer's so slow.

OpenStudy (trojanpoem):

Here , too.

OpenStudy (anonymous):

Which complex number does NOT fall within the shaded region in this graph?

OpenStudy (trojanpoem):

Choices ?

OpenStudy (trojanpoem):

There is plenty !

OpenStudy (anonymous):

-2i , 2-i , -2+3i , 1-i , -2-3i

OpenStudy (trojanpoem):

2-i

OpenStudy (anonymous):

Arrange the complex numbers in order according to the quadrant in which they appear, starting with the first quadrant. (3-4i),(-1-3i),(4+i),(-2+2i).

OpenStudy (trojanpoem):

Get the degree of each one can you ?

OpenStudy (trojanpoem):

Or easier way

OpenStudy (anonymous):

That's all it gives me. I like the new picture by the way.

OpenStudy (trojanpoem):

4 + i , -2 + 2i , -1 - 3 i , 3 - 4i

OpenStudy (trojanpoem):

x and y + in the first x ( _) and y + in the seconds - , - in 3rd and + , - in 4th

OpenStudy (trojanpoem):

Thanks.

OpenStudy (anonymous):

Arrange the complex numbers in order according to the quadrant in which they appear, starting with the first quadrant.

OpenStudy (trojanpoem):

I already arranged them

OpenStudy (anonymous):

NO problem and thank you so much Trojan.

OpenStudy (trojanpoem):

Is this another question ?

OpenStudy (anonymous):

In the graph, line( )passes through the complex number 1 + 2i and line( ) passes through -2 + 2i.

OpenStudy (trojanpoem):

1 + 4 i , 3 + 4 i , -2 + 2i , -2 + 4 i , - 1 -4 i , -11/5 -2 i , -3 -2 i , 3 -3 i

OpenStudy (trojanpoem):

E passes through 1 + 2 i

OpenStudy (trojanpoem):

G -> -2 + 2 i

OpenStudy (anonymous):

Which are real zeroes of this function? x^3 + 2x^2 – 9x – 18

OpenStudy (trojanpoem):

I can't see the number there is question marks in strange shape.

OpenStudy (anonymous):